Can I get detailed solution pls

If the size of a proton is of the order of 1 fm (10-15 m), what minimum kinetic energy must an electron beam have in order to “see” protons, inside a nucleus using electron scattering?

a. 10 MeV

b. 110 MeV

c. 1300 GeV

d. 1.2 GeV

e. none of the above
Answer #1

the\; formula\; of\; to\; calculate \; the\; energy \; is :

E=\frac{hc}{\lambda }

E=\frac{6.63\times 10^{-34}\times 3\times 10^{8}}{1\times 10^{-15}}
E=19.8\times 10^{-11} Joul E=19.8\times 10^{-11} \times 6.24\times 10^{9}GeV

\mathbf{E=1.2\; GeV}\; \; (Ans)